Consider making a mixture of 2,000 ml of A and B with 30 mole percent of A.

What volumes of pure A and B at 25℃ must be mixed to form the mixture also at 25℃?

(Partial molar volumes for A and B in the mixture of 30 mole percent A are 38 and 17 ml/mol and the molar volumes are 40 and 18 ml/mol.)
Answer #1

molar volume=V=x1+x2

V=0.3*38 + 0.7*17 = 23.3 ml/mol

Total volume=Vt=2000 ml

n = Vt/V = 2000/23.3 =85.8369 ml

n1=0.3*85.8369=25.751 mol

n2=0.7*85.8369=60.085 mol

V1t= 25.751*40 = 1030.04 ml

V2t = 60.085 * 18 = 1081.53 ml